Circular No. 8594

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005dp, 2005dq, 2005dr, 2005ds, 2005dt Further to IAUC 8592, E. Lee and W. Li report the LOSS discovery

of an apparent supernova (mag ∼ 18.5) on unfiltered KAIT images taken on Sept. 1.36 and 2.35 UT. SN 2005dt is located at α = 23^{h}23^{m}38^{s}.32, δ = −19^{o}00′24′′.5 (equinox 2000.0), which is 2′′.6 west and 11′′.7 north of the nucleus of MCG –03-59-6. KAIT images taken on 2004 Oct. 22.30 (limiting mag 19.5) and 2005 Aug. 14.31 (limiting mag 19.0) showed nothing at this position.

F. J. D. Serduke, M. R. Moore, W. Li, and A. V. Filippenko, University of California, Berkeley, report that inspection of CCD spectra (range 330–1000 nm), obtained on Sept. 2 UT with the Shane 3-m reflector at Lick Observatory, shows that SN 2005dp (IAUC 8591), SN 2005dq (IAUC 8592), SN 2005ds (IAUC 8592), and SN 2005dt (above) are young type-II supernovae, and that SN 2005dr (IAUC 8592) is a somewhat older type-IIn supernova. SN 2005ds may be of type IIn, as well, but higher-quality spectra are needed to confirm this.

SUPERNOVA 2005dl IN NGC 2276 S. Taubenberger and A. Pastorello, Max Planck Institut für Astro-

physik, Garching, on behalf of the European RTN collaboration (cf. IAUC 7987), report that a spectrum of SN 2005dl (cf. IAUC 8588), obtained on Aug. 30.02 UT by F. Hoyo with the Calar Alto 2.2-m telescope (+ CAFOS; range 340–880 nm), is that of a young type-II supernova. The noisy spectrum is characterized by broad P-Cyg features of the H-Balmer series and another P-Cyg line that can probably be attributed to He I 587.6nm. Adopting the NED recession velocity of 2410 km/s for the host galaxy (from Strauss et al. 1992, Ap.J. Suppl. 83, 29), the expansion velocity determined from Hβ is ∼ 9200 km/s. Supernova magnitudes estimated from broadband photometry obtained on Aug. 30.10 with the same instrument: B = 18.39 ± 0.20, V = 17.94 ± 0.15, R =17.35 ± 0.12, I =16.85 ± 0.10 (the errors take into account the uncertainties arising from a heavy contamination of the supernova light by the galaxy background). Host-galaxy extinction might be present, but the S/N of the spectrum is too low to unequivocally detect interstellar Na ID features.

NOVA IN THE SMALL MAGELLANIC CLOUD 2005 J. D. Neill reports that his observations reported on IAUC 8593 evi-

dently refer to Liller’s nova (cf. IAUC 8582).

2005 September 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green